OpenStudy (anonymous):

use the multiplication property of inequalities to solve the inequality. 1/8 x ≥ -5/6

hero (hero):

@Mimi_x3 can help you

hero (hero):

What the... really @Mimi_x3 You need wolframalpha for this?

OpenStudy (mimi_x3):

yeah..i dont have a calculuator

hero (hero):

You don't need a calculator for this?

OpenStudy (mimi_x3):

\[\frac{1}{8}x≥ -\frac{5}{6}\] \[x≥-\frac{5}{6}*8 =>x≥ \frac{-40}{6} \] yay!

OpenStudy (mimi_x3):

\[-\frac{40}{6} = -\frac{20}{3}\]